# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ 5.1.2 Econometric Evidence
FigureΒ 7 reports the results of the event study regression.151515The event study regression was not pre-specified and was suggested to us by seminar participants. Using administrative data on the date that a district first had access to STRV seed, we estimate the change in that districtβs EVI value before and after the introduction date. As we are looking at changes in district-wide EVI values, instead of changes in EVI values where flooding occurred, our event study is likely an underestimate of the effectiveness of STRVs in preventing flood damage. The event study supports our stated hypothesis. Prior to the introduction, EVI values were flat, fluctuating around zero change relative to the year before STRV availability (indicated as -1). In the years after STRVs were available in the district, EVI has increased relative to the value immediately before introduction. The longer the time period since introduction, the greater the increase in EVI. This is consistent with the slow but gradual increase in seed availability presented in FigureΒ 4. As the availability and amount of STRVs in the seed system has grown, so too has EVI values. This suggests that STRVs are effective in reducing yield loss due to flooding, as before their introduction flooded rice would have died, resulting in a lower EVI value than after STRV introduction.
Panel A of TableΒ 1 presents results from the DID specification. We include an indicator for if the EVI observation is from after 2009, meaning all districts are including in the βtreatedβ group starting in 2010, even if STRV seeds were not available until several years later. As can be seen in FigureΒ 4, this is the case for most districts, meaning our DID estimate will also likely underestimate the true effect. Each column in the table represents a regression with one of our four measures of flooding (cumulative, maximum, mean, or median). Our variable of interest is the interaction between the indicator for post-STRV introduction and the indicator for if the district is considered flood-prone using that flood measure of interest. The DID specification provides week support for our stated hypotheses. Point estimates across the four flood measures are consistently positive, though confidence intervals extend below zero. The point estimates on flooding are consistently negative and significantly different from zero. The only hypothesis that is not supported is that the post-STRV indicator would be zero. Absent flooding, growing STRVs should have no impact on yields, though in our results we see EVI is positive and significantly different from zero.
Panel B of TableΒ 1 presents results from the TWFE estimator. This is our preferred specification, as it incorporates both flood data and seed data from each district in each year. As with the DID results, each column in the table represents a regression with one of our four measures of flooding (cumulative, maximum, mean, or median). Results are similar in terms of sign and significance to the DID results. Flooding significantly decreases EVI while the availability of STRV seed has a very small but non-zero effect on EVI. The interaction between flooding and STRVs is also very close to zero, with confidence intervals centered around zero.
Summarizing the results, we find little evidence that the distribution of STRVs has had a significant impact on EVI, our proxy for yields, at a national-level. Only in the event study regression do we find statistically significant evidence for the positive impact of STRVs on EVI values in the years since STRVs became available in that district. We see marginally positive evidence in the DID specification, where flood-prone districts have higher EVI values after 2010 than before, though these estimates are not significant at standard levels of confidence. Finally, in the TWFE estimation, we see no significant evidence that EVI values are larger when it floods in a district with STRVs than EVI values when it floods in the same district before it had access to STRVs. Considering the preponderance of evidence, we conclude that the increased availability of STRVs has not reduced the loss of rice production due to flooding. This result stands in contrast to existing evidence from experimental data.
5.2 Household-Level Outcomes
5.2.1 Descriptive Evidence
We now turn to the households panel data to see if this more traditional data source supports or contradicts our EO-only approach for measuring long-term, large-scale impacts. First, we provide summary statistics for key variables across the three years of the panel (see TableΒ 2). The first three columns of the table report means and standard deviations for each variable in each year. The final three columns of the table report -values on year-to-year comparisons of means using a -test that allows for unequal variance. Statistically significant differences exist for almost every variable in almost every year-to-year comparison. Yields in 2014 and 2017 were around tons per hectare and increased to tons per hectare in 2022. Adoption was flat at eight percent of households for the first two years in the panel (in 2014, households adopted STRVs, in 2017, households adopted). Adoption grew substantially in the intervening five years, with ( households) adopting STRVs in 2022. Rice area has steadily declined over time and while statistically significant, the magnitude of the decline is not very large. In 2014, households had hectares in rice, which fell by a tenth of a hectare to in 2022. In terms of flooding, by most metrics households experienced the most flooding in 2014, followed by 2017, with the least flooding in 2022. However, in terms of the maximum intensity of flooding, 2022 was the worst. These differences highlight the importance of testing different flood metrics, as the duration (cumulative, mean, median), the extent (village), and the intensity (maximum) all are important considerations for the effectiveness of STRVs.
TableΒ 3 presents summary statistics by adoption status and for the data set overall. Households that adopt STRVs report higher yields in the years that they adopt compared to households not growing STRVs in a given year.161616Households can and do dis-adopt STRVs, though given how low adoption rates were in 2014 and 2017, dis-adoption is much less common that adoption. Recall also that we classify a household as adopting if they use STRVs on at least one of their rice plot. Thus, yields for adopters includes yields from both STRV and non-STRV seeds. See Online AppendixΒ E for plot-level analysis. The difference in yields between adopters and non-adopters suggests that the adoption decision is not random and may be correlated with household or farm unobservables, such as farmer ability or quality and location of plots. That said, we do not see significant differences in terms of rice area between adopters and not adopters. Surprisingly, we find that non-adopters experience great duration, extent, and intensity of flooding compared to adopters. One would assume households adopt because their plots are more prone to flooding. However, as adoption is not static or permanent, the difference in flood experience by adoption status may be the result of a non-adopting household experiencing an above average flood and then adopting STRVs the next year, when flood values are likely to regress to the mean. This potential time-varying adoption decision in response to flooding means that adoption is likely endogenous and also justifies our use of past flood experience as an IV for adoption.
5.2.2 Econometric Evidence
To understand the household-level impact of adoption, and to verify if they support our long-term, large-scale EO results, we implement two econometric approaches. First is a TWFE model that is the household equivalent to the district-level TWFE model. Second, given endogeneity concerns regarding the adoption decision, we estimate a TWFE-IV model in which the adoption decision is instrumented using lagged values of flooding, and their interactions, to calculate the probability that a household experiences flooding in each year of the panel data.
Panel A of TableΒ 4 presents results from the TWFE estimator which incorporates both contemporaneous household flood data and an indicator if the household was using STRVs on at least one plot. As with the district-level analysis, each column in the table represents a regression with a different flood metric (cumulative, maximum, mean, or median). To these, we also add a measure of flooding in the village, by calculating the village average of each householdβs maximum flood value, excluding the household in question. The TWFE results provide evidence that confirms the district-level EO analysis. Flooding significantly reduces yields and the use of STRVs during flooding has no significant impact on yields. The one results from the household data that does not match the EO data is the use of STRVs when it does not flood. We found a small positive impact in the EO data while we find no impact in the household data. This household result conforms to our hypothesis that STRVs should have no effect on yields absent floods.
Our final set of regressions (Panel B of TableΒ 4) uses historic flood data as an instrument for the adoption decision. The use of the IV addresses concerns that there may be unobserved transitory and idiosyncratic events that induce households to adopt but are not controlled for by the household and year fixed effects. Comparing the IV results to the TWFE results that do not employ the instruments, we observe more uncertainty in all of our estimates (wider confidence intervals), save for the estimates of the exogenous flood variable. This is consistent with the loss of efficiency in employing instrumental variables. With this loss of efficiency, it is unsurprisingly that we continue to find STRVs, absent flooding, have no significant impact on yields. For STRVβs impact on yields during flooding, coefficients are sometimes positive and significant, sometimes negative and significant, and sometimes not significant. Given this mixed evidence, we do not place a great degree of confidence in any one regression result, preferring to assume the mixed results are due to a loss of efficiency in estimation, which can increase both Type I and Type II errors.
The preponderance of evidence from the household-level analysis supports our conclusions from the district-level EO analysis. STRVs have a small or zero effect on yields when there is no flooding. Flooding, absent adoption, significantly decreases yields. Adopting STRVs and experiencing a flood has little to no effect on yields, relative to when households do not adopt and do not experience a flood. These results are all consistent with our EO analysis but are at odds with the existing body of evidence.
6 Failure to Replicate
The previous results lead to one obvious question: why do we fail to replicate the evidence in RCTs (Dar etΒ al., 2013; Emerick etΒ al., 2016) and field trials (Ismail, 2013) on the efficacy of STRVs? Two potential answers exist. First: there is no signal of STRV efficacy in the data because STRVs are not effective outside of a controlled experimental setting. We find this answer unlikely given the large number of observational and correlational studies that also find a positive association (Ismail etΒ al., 2013; Sanglestsawai etΒ al., 2014; Mishra etΒ al., 2015; Bairagi etΒ al., 2021) and the anecdotal reports in the press on the effectiveness of STRVs. That said, we cannot completely discount this answer, as there are no other long-term, large-scale causal studies in other countries to compare against. A second, and to our minds more plausible, answer is that measurement error in our data generates sufficient noise to obscure the true signal. We find this answer particularly compelling given that STRVs are a stochastic technology, dependent on floods occurring within a narrow window. As we are attempting to capture a higher order treatment effect, our analysis may be highly sensitive to mismeasurement and/or mis-classification.
We follow two avenues to explore the possibility that the true signal of STRV effectiveness has been lost in the noise inherent to conducting an impact evaluation in a data poor setting. First, we return to the flood data and work to make a number of improvements to capturing the length and depth of floods. Second, we use RCT data from Dar etΒ al. (2013) and Emerick etΒ al. (2016) to run Monte Carlo simulations on the sensitivity of their results to mismeasurement or mis-classification. In addition to the following exploratory analysis, we also conduct a number of pre-specified robustness checks. Summaries of those results, along with accompanying tables and figures, are in Online AppendixΒ D andΒ E. These robustness checks confirm that lack of statistically significant impact of STRVs on EVI or yields under a variety of variations to the data and econometric specifications.
6.1 Improved Flood Metrics
Two key challenges exist in measuring the impact of STRVs. The first is a paucity of data, which we address through our approach to reconstructing the past using EO data. The second is the βGoldilocks Problemβ or attenuation problem inherent in stress-tolerant seed varieties. For Sub1, any flooding outside the seven to 17 day window drives any impact on yields to zero. Additionally, inundation needs to be sufficiently deep to submerge the rice but not too deep (no greater than 25 cm). So, if our EO approach is to be successful, we need to ensure our flood maps capture the relevant flood event. Our main results are based on our pre-analysis plan, which in turn is based on what we knew was possible for capturing floods in the EO data. Subsequent to our initial analysis, we returned to the data and our fusion model in an attempt to refine the flood metrics. We generate new measures of flooding to more precisely estimate the depth and duration of the flood.
First, we create new pixel-level data where the value of the fractional index varies based on how we define the baseflow of water. To do this, we draw the distribution of flooding across the entire country for the entire year. We then define a pixel as being flooded if its fractional value is in the top five percent of values and zero otherwise. We repeat this process for every five percentile increment up to the percentile. The upper panel of FigureΒ 8 shows the results for all of Bangladesh. When the quantile is set at the top five percent of values, almost nowhere in Bangladesh would be considered flooded. When the quantile is set at the top 95 percent of values, almost everywhere in Bangladesh would be considered flooded. As we have no direct measure of the depth of water, varying the level of water that we consider baseflow allows us to indirectly capture flood depth. Lower quantiles require a higher threshold above baseflow (deeper water) to be considered a flood. Higher quantiles requires a lower threshold above baseflow (shallower water) to be considered a flood.
Second, we create three new flood metrics based on the new data that attempt to more directly capture floods that occur inside or outside the necessary βsweet spotβ for STRVs to have a measurable impact. These metrics are 1) the maximum number of consecutive flood days, 2) the number of days within the seven to 17 day window, and 3) a binary equal to one if the number of consecutive flood days fell within the window. We calculate each of these metrics at each quantile threshold. The lower panel of FigureΒ 8 graphs the distribution for the number of consecutive days of flooding for each quantile. At lower quantiles, where inundation needs to be extreme to be considered a flood, the duration of a flood lasts very few days. At higher quantiles, where inundation needs to be just slightly above baseflow to be considered a flood, the duration of flooding lasts almost the entire season.
Using the three new flood metrics measured at each of the 18 quantiles, we estimate TWFE models to determine the effect of STRVs on our four EVI measures during periods of flooding. This produces 216 different combinations of the data, which we represent using specification charts in FigureΒ 9. Each panel in the figure displays coefficient estimates and confidence intervals on the interaction of STRVs and a new flood metric for a specific EVI measure (cumulative, max, mean, median). Specifications are sorted based on coefficient size from smallest (left) to largest (right). The gray diamonds below the coefficients indicate which of the three measures of flood duration was used in the regression along with which quantile was used to determine depth or intensity of flooding. As one can see, the vast majority of results are statistically insignificant.
Yet there are patterns which are clearly visible in the results. A first cluster of outcomes in the upper left of the specification charts that represent negative and (sometimes) significant effects on EVI. These occur when we have very high thresholds for inundation to be considered a flood and when we focus on the seven to 17 day flood window. A high threshold means that inundation must be extreme to be considered a flood and only the most extreme events would fall within the window in which STRVs are effective. The coefficients are negative when using this data signals that setting such extreme thresholds are likely too high, and what we are capturing in reality are flood events beyond the capacity of STRVs. A second cluster is indicated by the downward sloping diamonds marking maximum number of days at decreasing quantile thresholds. Coefficients from regressions using these data are all statistically insignificant, with coefficient sizes centered around zero. This indicates that just counting the maximum number of consecutive days of flooding, without regarding to the window of STRV effectiveness, is not a useful measure of flooding at most quantile thresholds. Finally, there is a cluster of data in the center right of the charts that represent positive and (sometimes) significant effects on EVI. These results mostly occur when we use a flood indicator that accounts for STRVsβ window of effectiveness and when we set a threshold for flooding in the to percentile range.
There is a pessimistic and optimistic way to interpret the results in FigureΒ 8. A pessimistic interpretation is that, even when we increase the precision of our flood maps, STRVs have no significant impact on EVI. This could be because the technology is ineffective or because or EO methods remain unable to reconstruct the past with sufficient precision to pick up the STRV signal among the noise. This interpretation is pessimistic because it assumes either the method or the technology does not work. The pessimistic interpretation recognizes that there are some significant results for very specific flood measures, but discounts these as outcome occurring at random because we are running so many regressions. A correction for multiple hypothesis testing results in a loss of significance for all coefficients. The optimistic interpretation is that, given sufficient care in defining floods in the relevant way, STRVs have a significant impact on EVI. The problem is not the method or the technology but refining the method to a sufficient degree of precision so as to capture a treatment effect of a technology that only occurs under very specific circumstances. The optimistic interpretation discounts the plethora of null results as the outcome of mis-classification in what constitutes a flood and mismeasurement of the relevant flood window. The small number of positive and significant coefficients is not due to running many regressions but of defining flooding in the appropriate way. In the next section we turn to Monte Carlo simulations to try and assess the believability of the optimistic interpretation by determining just how sensitive results are to mis-classification and mismeasurement.
6.2 Monte Carlo Simulations
Evidence from experimental trials, both agronomic and economic, have reliably show positive effects of STRVs on yields during floods. One benefit of experimental data is that the experimenter typically has substantial control over how the experiment is implemented as well as how data is measured and collected. In trying to understand why we fail to replicate the existing experimental data, and to try and determine if the pessimistic or optimistic interpretation of our specification charts is more appropriate, we conduct a number of Monte Carlo simulations using the experimental data. Specifically, we use data that form the basis of Dar etΒ al. (2013) and Emerick etΒ al. (2016). We then, in separate Monte Carlo simulations, add noise to the flood data, the yield data, and the adoption indicator in order to simulate mismeasurement or mis-classification.
We begin by conducting a pure replication of the results in Dar etΒ al. (2013), which reports on the findings from the first year of an RCT in Odisha, India. They find consistent, positive and significant impacts on yields for adoption of Swarna-Sub1 during flooding, relative to both classic Swarna and other traditional varieties. This evidence is consistent with the field trial data presented in FigureΒ 1. We are also able to replicate these results using the complete two years of RCT data from Emerick etΒ al. (2016).171717Though the two papers rely on the same source of data, the regressions run in Dar etΒ al. (2013), which closely match tests for efficacy in the agronomic field trial literature, and not the specifications used in Emerick etΒ al. (2016). We are not replicating any specific result in Emerick etΒ al. (2016). Rather we are using the full two years of data to reproduce results that match Dar etΒ al. (2013) in magnitude, sign, and significance.
The simulations for mismeasurement in the yield and flood data are conducted in the same way but are completed separately. Here, we describe our method for injecting noise into the yield data but the method applies exactly to the flood data, as well. We start by calculating the mean and standard deviation of yield in the RCT data. We then draw, with replacement, from a normal distribution with mean zero and standard deviation equal to one percent of the standard deviation in the data . For each observation of yield, we add these randomly drawn values. We run the regression in Dar etΒ al. (2013) to estimate the impact of STRVs during flooding on the new yield data, which has the same mean but a larger variance. We repeat this 10,000 times. We then increase the noise by one percent, so that the new distribution that we draw from has mean zero and a standard deviation equal to two percent of the original standard deviation . We again rerun the regression for 10,000 draws from this new distribution. We conduct this process starting at one percent intervals from up to .
Results for both yield and flood are presented in the top two panels of FigureΒ 10. In this graph, we plot the distribution of -values on the interaction of STRV adoption and flood for each level of added noise (from to ). The vertical dashed red line marks . In the top column for both panels, all -values from the 10,000 simulations are statistically significant. This is because we are just replicating Dar etΒ al. (2013) and Emerick etΒ al. (2016) 10,000 times. When we add to the true yield data, of -values are still significant. But just adding two percent of the true standard deviation as noise to the yield data results in only of regressions producing significant results. Adding any more than nine percent of the true standard deviation to the true yield data, so that the yield data is distributed with mean and standard deviation results in only five to 10 percent of significant -values: the same amount of -values one would expect to be significant due to random chance with a or critical value. Results are similar for flooding, but the loss of significance happens even quicker, with only six percent of -values significant after adding just to the flood data.
We follow a similar process for simulating mis-classification in the adoption data. However, as adoption is a binary indicator, it is not useful to use the value of standard deviation to add noise to the data. Rather, we calculate the percentage of adopters in the data and then re-assign adopters and non-adopters at random. So, for one percent of the data we draw at random four households, determine if they were an adopter or non-adopter, and then switch their adoption status. We repeat this 10,000 times. As before, we start the process at zero, so that we replicate the original results, and then move upwards at one percent intervals until we are re-assigning half of all observations.
Results from the mis-classification simulations are graphed in the bottom panel of FigureΒ 10. Clearly results are much more robust to mis-classification of adoption status. Mis-classifying up to five percent of observations has no meaningful impact on the outcomes. Mis-classifying of outcomes results in about half of all regressions producing significant results. Only if one mis-classified of observations does the share of significant -values fall into the five to ten percent threshold, where their significance could be wholly ascribed to random chance.
The take-away from this exercise is that even results on the efficacy of STRVs using experimental data can be very sensitive to mis-measurement (though less so to mis-classification). Perturbing the true yield data by adding to each value a random number drawn from a distribution , or two percent of the standard deviation, produces non-significant results in nearly of regressions. Significance goes away adding even less noise to the flood data. In terms of what these means for our failure to replicate, the Monte Carlo simulations add support for the optimistic interpretation. STRVs are effective during floods and we can measure that effect using a geospatial impact evaluation that relies on EO data. But the method requires extreme care and precision in accurately capturing the variables of interest. The acute need for precision is likely the result of trying to measure higher order treatment effects from a stochastic technology.

### v2

@@ 5.1 Econometric Evidence
FigureΒ 8 reports the results of the event study regression. Using administrative data on the date that a district first had access to STRV seed, we estimate the change in that districtβs EVI value before and after the introduction date. As we are looking at changes in district-wide EVI values, instead of changes in EVI values where flooding occurred, our event study is likely an underestimate of the effectiveness of STRVs in preventing flood damage. Regardless of which EVI measure we use, the event study results generally support our stated hypothesis. Prior to the introduction, EVI values were flat, fluctuating around zero change relative to the year before STRV availability (indicated as -1). In the years after STRVs were available in the district, EVI tends to increase relative to the value immediately before introduction. This pattern is strongest when we use median EVI and weakest when we use the maximum value of EVI. Regardless of the EVI metric, the longer the time period since introduction, the greater the increase in EVI. This is consistent with the slow but gradual increase in seed availability presented in FigureΒ 7. As the availability and amount of STRVs in the seed system has grown, so too has EVI values. This suggests that STRVs are effective in reducing yield loss when flooding occurs, as before their introduction flooded rice would have died, resulting in a lower EVI value than after STRV introduction.
We present results for our DID model (and our TWFE model) in specification charts. These charts allow us to compactly summarize the size and significance of our coefficient of interest as we test more than 600 potential flood windows to determine if STRVs are effective during a Goldilocks flood. Recall from SectionΒ 2.1 that there is a range of potential flooding that could be considered βjust rightβ for STRVs to be effective. And recall from SectionΒ 3.3 that we generate a large set of potentially relevant flood windows by varying the start and end day of the flood window as well as the total length of the flood window. We then intersect these windows with the number of days of consecutive flooding measured at various quantiles above baseflow. Along the horizontal axis, model specifications are organized from smallest coefficient to largest. One can then read up the vertical axis using the gray markers to determine exactly what combination of flood window and flood quantile was used for that specific regression. Across the top are coefficient sizes with confidence intervals. Red bars signify negative and significant coefficients; black bars signify insignificant coefficients; and blue bars signify positive and significant coefficients. Given that we did not pre-specify these flood metrics, and in the interests of transparency, we report results from all regressions for each flood metric and for each EVI metric, not just those that conform to our hypotheses.
For our DID results, we use the various flood window-quantile combinations to determine median flooding in the data. When then classify a district as flood prone if it experiences flooding above the median level and its binary flood value is set to one. A district below the median value is classified as not flood prone and its binary flood value is set to zero. We include an indicator for if the EVI observation is from after 2009, meaning all districts are including in the βtreatedβ group starting in 2010, even if STRV seeds were not available until several years later. Our coefficient of interest is on the interaction of these two binary indicators . Panel A of FigureΒ 9 presents results from the DID specification with the cumulative EVI value as the dependent variable. We see that a large majority () of potential flood metrics produce negative or insignificant results. A small but not insubstantial set of flood metrics appear to be βjust rightβ in that when we use one of those flood metrics to define flood prone and interact it with the post-2009 indicator, it significantly increases cumulative EVI.121212Because the DID specification uses a binary indicator for flood prone, there ends up being almost no variation introduced by the different flood durations. If a district was classified as flood prone using the flood quantile at a window of five days, it was also classified as flood prone when the flood window was expanded to more days. This means that many of the 656 DID regressions we run are redundant. When we consider TWFE regressions, this is no longer the case.
That of 656 regressions turn our to be positive and significant at the -level clearly raises concerns that we are discovering significant results due to random chance. If STRVs are ineffective, meaning the true effect is zero, and we ran 100 regressions, we would expect a false positive in five of the regressions, given we have set our threshold for significance at . Across all panels in FigureΒ 9 we find a larger share of significant results than we would expect just by chance. There are also several regularities in our results that build confidence that they are not due to random chance. First and foremost is that fact that the significant results all come from flood measures at the to quantile. If results were truly random, we would expect some positive results at each quantile. Second, this consistent pattern is repeated across all four EVI measures. Again, if results were driven by random chance, then we would not expect consistency in which quantiles mattered as we changed how we measure the dependent variable.
For our TWFE estimation, we generate flood metrics in two different ways. The first is to use a dummy variable to indicate if flooding at a given quantile fell within a specific flood window. The second is to use the actual number of days that flooding at a given quantile was within a specific flood window. FigureΒ 10 presents results from the TWFE with the binary indicators for flooding. A pattern very similar that in our DID results is clearly visible. Focusing on Panel A, the majority () of potential flood metrics produce null or negative results. Of the 612 regressions, 90 produce positive and significant results.131313There are slightly fewer potential flood metrics than with the DID regressions because there are more empty sets at the intersection of quantile and flood window. These βjust rightβ floods occur when we consider flood measures produced at the to quantile, which substantially overlaps the quantiles that produced positive and significant results in the DID regressions. Again, these results are robust to using any of our four EVI measures. In terms of flood window, unlike in the DID regression, here a pattern begins to emerge. When the window is five to 10 days long, around of regressions produce positive and significant results. This falls to around of coefficients being significant when the flood window is longer than 10 days.
Turning to the TWFE results in FigureΒ 11, which uses the number of days in the window, we see the same pattern as previously in terms of quantiles that produce Goldilocks floods. And we see a stronger pattern in terms of which flood windows matter than previously evident. As before, results do not differ qualitatively when we use different EVI metrics, so we focus on results in Panel A. Of the 612 regressions, 98 () are significant, slightly larger than the share of positive and significant results in the DID and binary TWFE regressions. These positive and significant coefficients again come from the to quantile. In terms of which flood window matters, around of regressions are positive and significant when the window is between five and 10 days, with only around of regressions producing significant results when the window is greater than 10 days. When the window starts or ends, does not appear to matter much, with about of regressions producing positive and significant results regardless of whether the window starts at day five or day 15. This may suggest that, in terms of EO data at least, getting the correct intensity and duration of flooding is more important than getting correct the exact start and end day of the flood window (length).
Summarizing our results, we run 7,520 DID or TWFE regressions in which we interact different flood measures with proxies for adoption of STRVs. Across all specification and EVI measures, we find that of coefficients are positive and significant - many more than we would expect to find by chance if the true effect of STRVs were zero. Furthermore, the significant coefficients all come from floods categorized at the same quantiles above baseflow: between the and quantile. While not all flood windows at these flood quantiles produce positive and significant effects, no flood measured at a quantile above the or below the ever produces a positive and significant results. We take this as strong evidence that 1) STRVs have a positive and significant impact on rice production, as proxied by EVI, during flooding and 2) this effect only materializes during a Goldilocks flood. Both of these outcomes confirm the existing field trial and RCT evidence in that STRVs are effective in reducing yield loss to flooding and that the technology is stochastic, only becoming effective when flooding is βjust right.β
5.2 Robustness Checks
While our main results are consistent across specifications and changes to the dependent variable, there are a variety of concerns that other factors may be driving the results. Key among these are: 1) other abiotic stressors, 2) the level of aggregation, and 3) the use of EVI and administrative seed data as proxies for actual yield and adoption. In this section, we briefly summarize a series of checks on our results to determine if they are robust to these other factors.141414Figures and tables related to these results along with additional discussion and data considerations are in Online AppendixΒ E andΒ F.
First, our analysis might be picking up the impact of other abiotic stressors, as well as seed designed to mitigate those stresses, instead of the impact of STRVs. STRVs are just one example of a class of stress tolerant rice varieties, the next most important of which, in Bangladesh, is salinity tolerant rice. In coastal regions, flooding is often tidal, which carries with it (and leaves behind) salt. Potentially compounded this issue is work by Chen and Mueller (2018) that finds very little impact of flooding on rice yields, because farmers adjust to floods by delaying planting date, and instead finds significant impacts of soil salinity. To test if our results are picking up the impacts of salinity tolerant rice, and not submergence tolerant rice, we conduct our event study, DID, and TWFE analysis on our district-level data but exclude the 19 coastal districts where salinity can be a problem. Our event study results are more muted in their measured impact. But, our DID and TWFE tend to be stronger in that coefficients tend to be slightly larger and there are slightly more positive and significant outcomes. We take this as evidence that our results are driven by the effect of submergence tolerant rice on EVI during floods and not salinity tolerant rice.
Second, the choice regarding the unit of aggregation may be an important factor in determining our results. Our main results are at the district-level because that is the unit of observation for the administrative seed data. However, the district-scale measure of STRV seed availability is itself a noisy measure of actual adoption. Further, by aggregating flood and EVI, which is observed in pixels identified as rice, up to the district-level likely smooths over a lot of variation. To test if our results are driven by aggregation to the district-level, we conduct our event study, DID, and TWFE analysis at the upazila-level.151515Bangladesh is divided into 553 upazilas. We combine many extremely small upazilas in the city and metro area of Dhaka, where rice is not grown, to get a final dataset of 503 upazilas. Results from the event study are stronger in that EVI is larger and increases more rapidly after the introduction of STRVs than EVI does in the district-level panel. However, when we turn to our DID and TWFE results, we find very little evidence of significant impacts. In most of the specification charts, less than of regressions are significant, meaning we cannot discount that those significant results are purely due to chance; though the quantiles that do produce significant results are the same as those in the district-level analysis. One reason our non-event study results might not be robust to using upazila-level EO EVI and flood data is that the seed data is still at the district-level. This mismatch of units means that we are trying to explain intra-district variation in EVI using a single district-wide seed value. This type of mismatch can introduce noise into estimates and in fact we do find confidence intervals to be much larger in the upazila-level results. Coefficients of similar size in district-level and upazila-level specification charts have much larger confidence intervals in the upazila-level results than in the district-level results. However, this explanation is not totally satisfying since the upazila-level results are similar for both TWFE and DID regressions, the later of which do not use any seed data. Regardless of the explanation, what is clear is that our results tend not to be robust to disaggregating EVI and flood measures to the upazila.
Third, our results may be driven by our use of EVI as a proxy for yield and seed dissemination data as a proxy for adoption. Rice is not cultivated at the district-level, or event the upazila-level, but by individual households on small ( ha) fields. District-level EO measures may be noisy or biased proxies of the on-the-ground household reality. While no household data exists that allows us to perfectly test the robustness of our results, we do have the three-year Rice Monitoring Survey (RMS) panel that is representative of rice cultivation in Bangladesh, though only includes data post-STRV introduction. We generate the same set of flood metrics as before but using household GPS coordinates. We then estimate the TWFE model using measured household rice yield instead of EVI and household adoption status instead of rice seed availability.161616Online AppendixΒ F discusses limitations in the data both in terms of models we can estimate (event study and DID) and endogeneity of the adoption decision. Results are robust to the use of household-level data. Though only around of regressions produce positive and significant results, those results come from quantiles in the to quantile range. Unlike most of the district-level results, the flood window plays a much larger role in determining what exactly is a Goldilocks flood. Significant results only come from floods that fall in the five to ten day window. Floods longer than that never produce significant results. We take this as evidence that our results are not just the outcome of our use of aggregate proxies for yield, flooding, and adoption but that our EO results do in fact reflect the impacts of adopting STRVs by households in Bangladesh.